# On the Formalization of Cryptographic Migration

### v3

@@ Acknowledgments
Due to Lemma 2 the migration length $\ell(G)$ as a property of the migration graph $G$ is well-defined. It basically expresses how many independent migration steps one has to perform in order to fully migrate a system. Equivalently, we see that the set of components $V$ partitions into the migration clusters of its components, i. e. $V=\bigcup_{v\in V}c(v)$ and for all $v_{1},v_{2}\in V$, we either have $c(v_{1})=c(v_{2})$ or $c(v_{1})\cap c(v_{2})=\emptyset$. In other words the relation $v_{1}\sim_{c}v_{2}:\Leftrightarrow c(v_{1})=c(v_{2})$ is an equivalence relation. Informally, we have
$c_{9}$$c_{8}$$c_{7}$$c_{6}$$c_{5}$$c_{4}$$c_{2}$$c_{3}$$c_{1}$
Figure 3: The condensation graph $G/c$ for the migration graph from Fig. 2.
@@
For a migration graph $G=(V,E)$ the graph $G/c$ is acyclic.
Proof.
Assume $G/c$ is not acyclic, that is, there are $v_{1},v_{2}\in V$ such that $c(v_{1})\to^{+}c(v_{2})\in(E/c)^{+}$ and $c(v_{2})\to^{+}c(v_{1})\in(E/c)^{+}$ as usual denoting by the superscript $+$ the irreflexive transitive closure of the dependency relation. But then, by definition of $c$, we have $c(v_{1})=c(v_{2})$. ∎
@@
Figure 4: The migration graph from Fig. 2 with its migration clusters and canonical migration times. In the example with $12$ components, we have $\ell(G)=9$ and $d(G)=5$.
[…]
@@
-D Simple Trade-Offs
First of all, there is a simple trade-off between the number of migration clusters in a graph $G$ and the size of the clusters:
@@
Assume that for a migration graph $G=(V,E)$ the migration length is bounded by a constant $k$, i. e. $\ell(G)\leq k$. Then, there is a component $v\in V$ with #$\#c(v)\geq n/k$.
@@
By definition $\ell(G)=\#(V/c)$ and $n=\sum_{c(u)\in V/c}\#c(u)$. Let $v$ be a component with a maximally large migration cluster, i. e. $v=\mathop{\text{argmax}}\nolimits_{u\in V}\#c(u)$. In general $v$ is not uniquely defined, but for our existential statement, this does not trouble us: Namely, for such $v\in V$, we have $n=\sum_{c(u)\in V/c}\#c(u)\leq\sum_{c(u)\in V/c}\#c(v)=\ell(G)\cdot\#c(v)\leq k\cdot\#c(v)$ and the claim follows. ∎
Also locally, there is a similar trade-off between the size of clusters and the number of independent components:
@@
Consider a component $v$ in a migration graph $G$ with $m\in\mathbb{N}_{>0}$ direct dependencies. Assume there are at most $k$ independent clusters within those $m$ components. Then there is a migration cluster of size $s$ with $s\geq m/k$. Conversely, if each migration cluster has size at most $s$ then we have $k\geq m/s$ migration clusters.
@@
Assume each migration cluster has size $s<m/k$. Since there are at most $k$ independent clusters, we have at most $k\cdot s<m$ components. Contradiction. ∎
Thus, we see that we have to expect either many independent components or at least one large migration cluster. Another simple trade-off emerges, when we bound the number of descendants of any component in the graph suitably. Then, we get a large migration length as seen from
Theorem 4.
Let ${f\colon\mathbb{R}\rightarrow\mathbb{R}}$ be a smooth function in a real variable with $f\in o(n)$, e. g. $f(n)=\mathop{\text{log}}\nolimits n$. Assume for any component $v$ in a migration graph $G=(V,E)$ with $n$ components, we have for each $k\geq 0$ the bound #$\#(v\to^{k}V)\leq f(n)$. Then $\ell(G)\geq n/f(n)$. In other words, the number of elements in the set of components with a common ancestor of distance $k$ has a trade-off with the length of the graph.
Proof.

### v4

@@ III-D Migration Strategies
For a migration graph $G=(V,E)$ the graph $G/c$ is acyclic.∎
@@
The overall structure of the migration graph $G$ is thus as follows (cf. Fig. 3):
@@
$v_{10}$$v_{11}$$v_{12}$$v_{9}$$v_{8}$$v_{7}$$v_{6}$$v_{5}$$v_{4}$$v_{2}$$v_{3}$$v_{1}$ $c_{1}$ $c_{2}$ $c_{3}$ $c_{4}$ $c_{5}$ $c_{6}$ $c_{9}$ $c_{8}$ $c_{7}$ $T(1)$ $T(2)$ $T(3)$ $T(4)$ $T(5)$
Figure 3: The migration graph from Fig. 1 with its migration clusters (denoted by dashed boxes) and canonical migration times. In this example with 121212 components, we have $\ell(G)=9$ and $d(G)=5$.
IV Assessing the Difficulty of the Migration Problem
We now show that, as expected, any sufficiently large migration project splits into many dependent migration steps that cannot be parallelized and provide a formal proof of this statement.
From classical results in combinatorial analysis we know the following about random decompositions of the migration graph into its clusters.
IV-A Random Migration Clusters
Since every migration graph decomposes disjointly into its migration clusters, we now study the expected properties of these migration clusters. Let $n$ be the number of components. There are $B_{n}$ different possible partitions of $n$ components in disjoint subsets, where
$B_{n}=\sum_{k=0}^{n}{n\choose k}B_{k}=\frac{1}{e}\sum_{k=0}^{\infty}\frac{k^{n}}{k!}$
is the $n$th Bell number. More precisely, there are $n\brace k$ partitions of our components in $k$ distinct non-empty subsets, where $n\brace k$ are called the Stirling numbers of the second kind. Obviously, we have $B_{n}=\sum_{1\leq k\leq n}{n\brace k}$.
IV-A1 Sampling Migration Clusters
Consider now a migration project with $m$ different sub-projects and suppose each of our $n$ components is assigned to one of the $m$ sub-projects uniformly at random. Let $\Pi_{nm}$ be the random variable defining the resulting partition $\pi$ of our components. We have:
$\mathop{\text{prob}}\nolimits(\Pi_{nm}=\pi)=\frac{m^{\underline{k}}}{m^{n}}$
if #$\#\pi=k$, where $m^{\underline{k}}=m\cdot(m-1)\cdots(m-k+1)$ is the falling factorial function. In other words: if there are $k$ different sub-projects with at least one component, then the probability of creating one such partition is given by the formula above. For the distribution of the number of non-empty different sub-projects we have
$\mathop{\text{prob}}\nolimits(\#\pi=k)={n\brace k}\frac{m^{\underline{k}}}{m^{n}}.$
Note that for arbitrary migration graphs, we do not know the distribution of the different partitions. We have the following
Theorem 1 (Proposition 2 in [25]).
Let $M$ be the random variable with values in $\mathbb{N}_{\geq 1}$ and suppose given $M=m$ that $n$ balls labeled by $\mathbb{N}_{\leq n}$ are thrown independently and uniformly at random into $m$ boxes. Then the following two statements are equivalent:
  1. 1.
    $M$ is distributed via $\mathop{\text{prob}}\nolimits(M=m)=m^{n}p_{m}/B_{n}$ for $m\in\mathbb{N}_{\geq 1}$, where $p_{m}$ is a probability distribution on $\mathbb{N}_{\geq 0}$ whose first $n$ factorial moments are identically equal to 111. One may for example take $p_{m}=1/(em!)$.
  2. 2.
    $\Pi_{nM}$ has uniform distribution over the set of all partitions of $n$ components.
The Theorem gives a constructive way of uniformly generating a migration cluster: First sample $m\in\mathbb{N}_{\geq 1}$ following the above distribution, then distribute the $n$ components uniformly in $m$ bins. The resulting partition $\pi$ will be a uniformly selected partition of our $n$ components. We will now study the expected properties of the outcome.
IV-A2 Distribution of Random Migration Clusters
Studying the distribution of random partitions of sets with $n$ elements is a classical problem. Set into the context of migration problems, we state
Theorem 2 (Summary of [26] and Chapter 4, Theorem 1.1 in [27]).
The number of migration clusters $M$ is expectedly $\mathsf{E}(M)\sim n/\mathop{\text{log}}\nolimits n$ and the expected size of each migration cluster is $\mathop{\text{log}}\nolimits n$. Also, almost all partitions have approximately $e\mathop{\text{log}}\nolimits n$ distinct cluster sizes. More specifically the distribution of the number of migration clusters $M$ converges asymptotically to a normal distribution with mean $\mu=n/\mathop{\text{log}}\nolimits n$ and standard deviation $\sigma=\sqrt{n}/\mathop{\text{log}}\nolimits n$ for $n\to\infty$.
Furthermore, the expected maximum size of a migration cluster is, for growing $n$, asymptotic to $er-\mathop{\text{log}}\nolimits\sqrt{r}$, where $r\cdot\mathop{\text{exp}}\nolimits(r)=n$ and $r=\mathop{\text{log}}\nolimits n-(1+o(1))\mathop{\text{log}}\nolimits\mathop{\text{log}}\nolimits n\sim\mathop{\text{log}}\nolimits n$ [28]. This means that we typically expect lots of small clusters and we do not expect exceptionally large ones. Results on the speed of convergence are beyond the scope of this article. We now compute an upper bound on the expected number of migration clusters when looking independently at $s$ randomly selected migration projects.
@@ IV-A2 Distribution of Random Migration Clusters
Let $M_{1},\ldots,M_{s}$ be the resulting numbers of migration clusters and define $X=\mathop{\text{max}}\nolimits_{1\leq i\leq s}M_{i}-\mu$. Then conveniently $\mathsf{E}(X)=0$. Since for a fixed real constant $t$ the function $x\mapsto\mathop{\text{exp}}\nolimits(tx)$ is convex, by Jensen’s inequality we have:
$\leq\sum_{i=1}^{s}\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))=s\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$
Here, the last equality follows from the fact that the moment generating function $\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$ for a normal distribution with mean zero is equal to $\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$ and in particular independent of the index-variable $i$. Rewriting this, $\mathsf{E}(X)\leq\frac{\mathop{\text{log}}\nolimits s}{t}+\frac{t\sigma^{2}}{2}$. This bound is minimal at $\frac{\sqrt{2\mathop{\text{log}}\nolimits s}}{\sigma}$ since its derivative in direction $t$ is $\frac{\sigma^{2}}{2}-\frac{\mathop{\text{log}}\nolimits s}{t^{2}}$. Thus, $\mathsf{E}(X)\leq\sigma\sqrt{2\mathop{\text{log}}\nolimits s}$. ∎
In practice, the Lemma implies that we would have to sample exponentially many migration graphs (in the number of components $n$) in order to expectedly observe graphs with asymptotically more than $n/\mathop{\text{log}}\nolimits n$ migration clusters. Thus, the statement from Theorem 2 applies in practice all the time, at least for sufficiently large $n$. It remains to study the dependencies between the migration clusters, i. e., the structure of the condensation modulo the migration clusters.
TABLE I: Numerical evaluation of the expected behavior of migration clusters for migration projects with $n$ components rounded to the nearest integer. We see that one does not expect large mutual dependencies but the expected number of independent clusters is vast and the chain-dependencies are also large when we bound the number of migration clusters with fixed canonical migration time by $w(n)=\mathop{\text{log}}\nolimits n$ or $w(n)=\sqrt{n}/\mathop{\text{log}}\nolimits n$, respectively.
[…]
IV-B Properties of the Condensation Graph
According to Theorem 2, we consider a condensation graph, i. e., the directed acyclic graph resulting from the migration graph on $n$ components modulo its strongly connected components, with expectedly $n^{\prime}=n/\mathop{\text{log}}\nolimits n$ vertices. Since we are considering a single migration project, we can assume that this graph is connected. How can we model this graph?
A first idea would be to use the standard Erdős–Rényi model [29], where the graph is constructed by connecting the components independently at random with a fixed probability $p$. One can ensure that the result is acyclic by defining any linear order on the vertices and include a directed edge only if it obeys the selected ordering. Each such edge is included in the graph with probability $p$, independently from every other edge. For such graphs, we have
Theorem 4 (Theorem 4.1 in [30], a variant of Theorem 1 in [31]).
Let $G_{n,p_{n}}$ be a random $n$-vertex graph, where each arc is present with probability $p_{n}=\frac{\mathop{\text{log}}\nolimits n+c_{n}}{n}$. Then, for growing $n$, the probability that $G$ is connected is given by
$\lim_{n\to\infty}\mathop{\text{prob}}\nolimits(G_{n,p_{n}}\text{ connected})=\begin{cases}0&\text{, if $c_{n}\to-\infty$}\\ e^{-e^{-c}}&\text{, if $c_{n}\to c$ const.}\\ 1&\text{, if $c_{n}\to\infty$}\\ \end{cases}$
Note that we can directly apply this Theorem to our directed acyclic graph situation, since such a graph is connected if and only if the undirected variant of it is connected.
As our goal is to prove a lower bound on the number of edges in the condensation graph (with $n^{\prime}$ vertices!) in our model, we assume now the sparsest case $c_{n^{\prime}}\to c$ constant for which the graph is connected with high probability. Note that the doubly-exponential term in Theorem 4 approaches 111 already for comparatively small $c$ very quickly. For example, already for $c=10$, the value of $e^{-e^{-c}}\geq 0.99995$. But also for any $c>0$, we obtain for the number of edges in the condensation graph expectedly asymptotically $\frac{\mathop{\text{log}}\nolimits n^{\prime}+c_{n^{\prime}}}{n^{\prime}}\cdot\frac{n^{\prime}\cdot(n^{\prime}-1)}{2}=\frac{1}{2}((n^{\prime}-1)(\mathop{\text{log}}\nolimits n^{\prime}+c_{n^{\prime}}))\sim\frac{1}{2}n^{\prime}\mathop{\text{log}}\nolimits n^{\prime}$. That means that even in the sparsest case, we expect a considerable number of dependencies between the migration clusters.
We now analyze what happens if we bound the number of clusters with a common ancestor of distance $k$ (the maximum being the width $w$ of the graph), i. e., #$\#(c(v)\to^{k}V/c)$ by a small function $w(n^{\prime})=o(n^{\prime})$ in the size $n^{\prime}$ of the condensation graph $G/c$.
@@ IV-B Properties of the Condensation Graph
[…]